OpenStudy (anonymous):

In right ∆ABC, m B = 90°, m C = 40°, and BC = 10. What are the other two side lengths of the triangle?

OpenStudy (anonymous):

AC = 13.1, AB = 8.4 AC = 18.7, AB = 15.8 AC = 14.3, AB = 10.2 AC = 15.6, AB = 11.9

OpenStudy (anonymous):



OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

\[\frac{ AC }{ \sin90 }=\frac{ BC }{ \sin50 }\] BC=10

OpenStudy (anonymous):

can you calculate AC?

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

how? AC=\[\frac{ 10 }{ \sin50 }*\sin90\] use a calculator

OpenStudy (anonymous):

Not understanding this either.